### v1

# Real-Time Wayfinding Assistant for Blind and Low-Vision Users

Abstract.
Navigating unfamiliar places continues to be one of the most persistent and essential everyday obstacles for those who are blind or have limited vision (BLV). Existing assistive technologies, such as GPS-based navigation systems, AI-powered smart glasses, and sonar-equipped canes, often face limitations in real-time obstacle avoidance, precise localization, and adaptability to dynamic surroundings. To investigate potential solutions, we introduced PathFinder, a novel map-less navigation system that explores different models for understanding 2D images, including Vision Language Models (VLMs), Large Language Models (LLMs), and employs monocular depth estimation for free-path detection. Our approach integrates a Depth-First Search (DFS) algorithm on depth images to determine the longest obstacle-free path, ensuring optimal route selection while maintaining computational efficiency. We conducted comparative evaluations against existing AI-powered navigation methods and performed a usability study with BLV participants. The results demonstrate that PathFinder achieves a favorable balance between accuracy, computational efficiency, and real-time responsiveness. Notably, it reduces mean absolute error (MAE) and improves decision-making speed in outdoor navigation compared to AI-based alternatives. Participant feedback emphasizes the system’s usability and effectiveness in outside situations, but also identifies issues in complicated indoor locations and low-light conditions. Usability testing revealed that 73% of participants understood how to use the app in about a minute, and 80% praised its balance of accuracy, quick response, and overall convenience. By addressing critical usability and performance criteria through a low-cost, efficient, and scalable design, our study helps to advance AI-driven assistive technology for people with BLV, supporting greater independence and accessibility in daily navigation.
Blind and low vision (BLV) people, Indoor and outdoor navigation, Assistive technology, Large Language Models (LLMs), Depth estimation, Depth First Search (DFS), Computer Vision
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing Accessibility technologiesccs: Computing methodologies Computer vision problemsccs: Information systems Navigation-based services [^1] Code and dataset are available at: https://github.com/dabbrata/AI_Assisted_Nav_DFS and https://www.kaggle.com/datasets/artemmmtry/diode-a-dense-indoor-and-outdoor-depth-dataset
1. Introduction
Navigating unfamiliar environments is a significant challenge for blind and low-vision (BLV) individuals, both indoors and outdoors. While traditional mobility aids such as white canes and guide dogs provide essential support, they have limitations in detecting obstacles and identifying clear paths in complex environments. To enhance mobility, BLV individuals rely on assistive technologies such as GPS-based navigation systems, AI-powered smart glasses (Waisberg et al., 2024), and sonar-equipped smart canes (Messaoudi et al., 2020). These solutions provide audio or haptic feedback for obstacle detection and route guidance. However, despite these advancements, existing technologies still face limitations in real-time navigation, precise obstacle avoidance, and environmental awareness, particularly in dynamic and unfamiliar scenarios.
Many GPS-based assistive technologies, including Be My Eyes (Redaktion, 2023), Seeing AI (Microsoft, 2017), and Aira (Aira Tech Corp, 2015), offer voice guidance and real-time assistance. However, they often struggle with providing precise localization, real-time obstacle avoidance, and effective pathfinding in unstructured environments. Additionally, AI-powered solutions, such as computer vision-based smart glasses (AlSaid et al., 2019), require high computational power, making them costly and less accessible for many users. Given these challenges, there is a need for a more efficient, low-cost, and user-friendly navigation solution that leverages image processing and depth estimation techniques to assist BLV individuals effectively.
In this paper, we examine various navigation techniques, including those that use Vision Language Models (VLMs), and then present a computer vision-based algorithm for recognizing free space and enabling pathfinding. The system also incorporates monocular depth estimation to generate depth maps, which, when processed using LLMs, improve obstacle detection and free path identification. Finally, we propose a novel image-processing approach that leverages Depth-First Search (DFS) on depth images to identify the longest free path, ensuring optimal route selection for BLV individuals. This computationally efficient technique enhances real-time navigation by providing accurate obstacle-free paths while maintaining low processing costs. Compared to existing AI-powered solutions, our method offers a user-friendly, computationally efficient, and cost-effective alternative without compromising accuracy or responsiveness, making it a practical and accessible solution for BLV individuals to navigate safely and independently.
To evaluate the effectiveness of our system, we aim to answer the following research questions:
  • RQ1: How useful is our proposed navigation scheme for BLV people?
  • RQ2: How effective are AI-powered solutions for real-time navigation compared to our method?
  • RQ3: How well does our method address the shortcomings of AI-powered solutions for real-time navigation?
  • RQ4: How well does the implemented system address BLV individuals’ needs, and what are their impressions of its effectiveness?
To answer RQ1, we conducted preliminary studies on BLV individuals’ navigation habits and identified key challenges, such as difficulties in anticipating environmental layouts, avoiding obstacles, and maintaining a straight path in dynamic settings. Our approach aims to overcome these challenges through a cost-effective, real-time navigation aid that can be used with a single-camera smartphone.
To answer RQ2, we compare our method with existing AI-powered solutions that rely on high-end computational models for navigation. While AI-based systems provide valuable insights, they often require significant resources and may not perform optimally in all environments. Our approach, utilizing depth maps and DFS-based free-path detection, offers a lightweight yet efficient alternative.
To answer RQ3, we analyze the primary limitations of current AI-powered navigation solutions, including high computational costs, reliance on expensive hardware, and potential inaccuracies in real-time obstacle detection. Our method addresses these issues by using a more accessible and computationally efficient approach, ensuring better accuracy, faster processing, and enhanced usability.
To answer RQ4, we are conducting user studies and surveys in February 2025 to evaluate how well the implemented system addresses BLV individuals’ needs and their impressions of its effectiveness. This evaluation aligns with ongoing advancements in assistive technology, urban infrastructure developments, and relevant policy changes. Early surveys assessing public interest in assistive navigation solutions have highlighted key concerns regarding privacy, data security, and long-term usability (Martin et al., 2010). Our study considers these factors to refine the system, enhance user trust, and improve adoption, ensuring it effectively meets the real-world navigation challenges faced by BLV individuals.
While our system offers significant advantages in terms of accessibility, efficiency, and accuracy, we acknowledge potential limitations related to privacy concerns and adaptability in highly dynamic environments. Our research adds to the larger goal of increasing accessibility and independence by creating a user-friendly, cost-effective, and efficient navigation system for BLV persons. Furthermore, the approaches investigated in this study can contribute to future advancements in real-time assistive technology for people with visual impairments, as well as inspire breakthroughs in autonomous navigation systems for personal and public use.
2. Related Works
Developing successful navigation systems for blind and low-vision (BLV) people has been a major focus of assistive technology research. Various methodologies have been investigated, ranging from GPS-based solutions and depth-sensing techniques to AI-powered systems and infrastructure-free navigation tools. While many of these technologies strive to increase mobility, obstacle detection, and spatial awareness, they frequently confront hurdles in real-time processing, environmental adaptability, and user accessibility. This section examines existing research on BLV navigation systems, depth estimate methods, and AI-powered assistive technologies, emphasizing their contributions, limits, and implications for the development of our proposed solution. Table 1 highlights significant studies on BLV navigation and accessibility, including techniques, conclusions, and limitations.
2.1. Navigation Systems for BLV Individuals
Navigating unknown regions poses significant obstacles for blind and low-vision (BLV) people, demanding intensive study into assisted navigation technologies. A variety of approaches have been investigated, ranging from GPS-based solutions to sensor-driven indoor navigation systems.
Traditional GPS-based technology have been important in assisting BLV patients. Microsoft Soundscape (Rickert, 2010), for example, improves spatial awareness with 3D audio cues. Liu et al. (Liu et al., 2022) studied user interaction with Soundscape and discovered that spatial audio callouts considerably improved usability and retention. Their investigation did, however, identify limits in exact localization and real-time obstacle avoidance, both of which are significant challenges in GPS-based systems.
To address the limitations of GPS-based technology, other solutions based on ambient sensors and computer vision have been investigated. Jain et al. (Jain et al., 2023) describes a system that uses current street cameras to deliver real-time aural feedback to BLV pedestrians. By using computer vision algorithms to street camera feeds, the system provides users with detailed spatial information that aids in obstacle avoidance and safe street crossing. While this strategy improves real-time navigation, its reliance on public monitoring infrastructure limits its scalability.
For indoor navigation, Tsai et al. (Tsai et al., 2024) proposed inertial-based iOS applications—Wayfinding and Backtracking—that use a smartphone’s internal sensors for pathfinding without relying on external infrastructure like Bluetooth beacons. The Wayfinding software enables hands-free navigation along established routes, while Backtracking allows users to retrace their steps in unfamiliar areas. Although these systems are useful, they suffer from localization drift—a progressive rise in positioning errors over time—which leads to lower navigation accuracy, especially in bigger or featureless areas.
Depth sensing combined into portable devices has also emerged as a viable solution. See et al. (See et al., 2022) created a smartphone application that uses depth cameras to identify obstacles and recognize objects. This infrastructure-free technique provides real-time environmental awareness via a mobile device. In controlled experiments, the system detected obstacles accurately, although performance varied depending on environmental conditions such as lighting and item size.
Robotic help expands the capabilities of navigation aids. Kuribayashi et al. (Kuribayashi et al., 2023) built a suitcase-shaped robot (PathFinder) that helps blind people navigate unfamiliar indoor places by dynamically recognizing junctions and reading signs. The solution gives real-time input without relying on prebuilt maps, giving users more flexibility. However, it needs users to recall routes while navigating, which might raise cognitive burden and impair usefulness.
Ren et al. (Ren et al., 2023) proposed RouteNav, which is a hybrid system that combines GPS localization and inertial dead reckoning to increase navigation accuracy in complex areas like transportation hubs. This system mitigates the impact of GPS inaccuracy by segmenting environments into a tile-based navigation system that provides multimodal feedback such as voice, haptics, and aural signals. While the system is excellent in organized surroundings, it still faces issues with localization drift and GPS signal discrepancies, particularly in underground or open locations.
Table 1. Summary of Related Works on BLV Navigation and Accessibility
Author(s) Focus Area Methodology Key Findings Limitations
Liu et al. [6] GPS-based outdoor navigation Uses spatial audio callouts for location awareness Improves spatial perception and usability Limited accuracy in localization and obstacle avoidance
Tsai et al. [11] Indoor navigation without external infrastructure Uses smartphone sensors for path tracking Enables retracing steps and predefined routes Suffers from localization drift over time
See et al. [8] Depth camera-based obstacle detection Uses embedded depth cameras for feedback Provides auditory/haptic feedback for real-time navigation Performance varies by environment, limited object categories (e.g., furniture, doorways)
Seo et al. [9] AI-driven data visualization for BLV users Combines text, audio, and haptic feedback Enhances structured interaction with complex data AI hallucinations, difficulty in prompt engineering
Zhang et al. [12] Accessible 2D data visualization for BLV users Multimodal interfaces (touch, audio, haptic) Supports scatter plots, bar charts, and line charts Complex gestures, small sample sizes in testing
Jain et al. [9] Real-time urban navigation using public cameras Analyzes live video feeds for path guidance Adapts to dynamic obstacles in cities Requires permanent infrastructure, limiting scalability
Shaheen [10] Scalability in educational accessibility Assesses assistive tech usage in education Highlights lack of BLV-friendly resources Requires external assistive devices for basic access
2.2. AI-Powered Scene Description and Accessibility Tools
Beyond navigation, AI-powered scene description tools have played an important role in increasing BLV users’ environmental awareness. These applications combine computer vision and natural language processing to deliver real-time descriptions of the environment, objects, and people in order to promote user independence. However, issues such as trust in AI outputs, description accuracy, and user experience remain major concerns.
Penuela et al. (Gonzalez Penuela et al., 2024) looked into how BLV users engage with AI-powered scene description apps, including their trust, satisfaction, and usage habits. The study discovered that users regularly used these tools for object recognition and scene assessment in familiar interior settings, based on a two-week diary study with 16 participants. Interestingly, despite occasional mistakes in AI-generated descriptions, users frequently relied on the system, demonstrating a nuanced link between trust and perceived utility. The study also stressed the significance of increasing AI accuracy and user customisation choices in order to improve satisfaction and usability.
Another interesting method (See et al., 2022) is to include depth photography into assistive applications to improve obstacle identification and spatial awareness (See et al., 2022). A depth camera-based navigation tool was created to help users recognize barriers and locate things by providing aural and haptic input, removing the need for extra external hardware. The technology successfully supplied real-time input but had limitations in different contexts. Furthermore, the object detection model was limited to specific categories, such as furniture, entrances, and everyday domestic things, restricting its applicability in unknown environments.
To broaden the scope of AI-powered accessibility, Seo et al. (Seo et al., 2024) explored how BLV users interact with multimodal large language models (LLMs) to interpret complicated data visualizations. The study launched maidrAI, an AI-powered data visualization tool that combines text, audio, and haptic input, allowing BLV users to interact with data using organized, customized descriptions. User research found that participants preferred clear, organized outputs and valued the opportunity to tailor AI-generated solutions. However, issues such as AI hallucinations (inaccurate or falsified information) and the difficulty of rapid engineering remained, emphasizing the need for more dependable and user-friendly AI systems.
Efforts to increase data accessibility for BLV users, ChartA11y(Zhang et al., 2024) allows blind people to interact with 2D data visualizations via multimodal interfaces on their smartphones. The technology, which combines touch-based exploration with audio and haptic feedback, gives users access to scatter plots, bar charts, and line charts. User tests demonstrated the usefulness of mixing structured navigation with direct touch interactions, however obstacles such as gesture complexity and small sample sizes in testing indicated areas for additional research.
2.3. User Engagement and Broader Assistive Technologies
User-centered design and engagement strategies are crucial for developing assistive technologies that meet the requirements and preferences of blind and low-vision (BLV) people. Several studies have looked into how design decisions affect user adoption, satisfaction, and long-term involvement with navigation and accessibility technologies.
In their study, Liu et al. (Liu et al., 2022) examined user behavior using Microsoft Soundscape to better understand the aspects that influence long-term engagement. The study found that early user interactions, particularly those during the first week, were major indicators of continuing use. Features such as spatial audio callouts and street previews helped to retain users. However, the study also noted difficulties in extrapolating findings to other navigation tools and advocated for the incorporation of qualitative user feedback to better customize future solutions.
Tsai et al. (Tsai et al., 2024) highlight the need for flexible, infrastructure-free navigation tools by introducing iOS applications for interior navigation using a phone’s built-in inertial sensors. By eliminating the requirement for external beacons or pre-installed infrastructure, the system enabled hands-free navigation, which is critical for BLV people who frequently rely on canes or guide dogs. While user studies validated the system’s usefulness, concerns like as localization drift and decreased accuracy in wide, open settings hampered its general applicability.
Educational accessibility for BLV students has also received substantial attention in assistive technology research. Shaheen (2024) (Shaheen, 2024) investigated the difficulties encountered by BLV students in K-12 school, particularly in using instructional tools. The study discovered that many children gained assistive technology (AT) literacy outside of school due to inadequacies in formal education assistance. Complex user interfaces, unlabeled buttons, and irregular navigation frameworks hampered academic involvement, particularly in STEM disciplines.
Urban navigation provides new issues, notably in picking the most relevant information to assist BLV users during mobility activities. Hoogsteen et al. (Hoogsteen et al., 2022) explored which environmental characteristics are easiest to describe for blind pedestrians. Intersections, traffic lights, and crosswalks were judged to be the most useful, while things such as poles or fire hydrants were deemed less relevant. Interestingly, the study found that early and late blind people had varied information needs, with late blind users requiring more extensive environmental descriptions. The study provided a structured framework for selecting information in assistive navigation systems, with the goal of lowering cognitive burden and improving spatial awareness.
2.4. Gaps in Existing Research
Despite major advances in assisted navigation and AI-powered accessibility technologies, issues such as processing efficiency, real-time flexibility, usability, and scalability continue to impede their practical adoption by blind and low-vision (BLV) people.
One significant barrier is the processing cost of AI-powered systems. Many depth estimation and object identification methods need a lot of computing power, which makes them unsuitable for mobile deployment. While some technologies solve these restrictions, relying on specialist hardware like depth cameras and LiDAR raises costs and reduces accessibility (See et al., 2022). Although lightweight AI models have been proposed, they frequently forfeit accuracy, limiting their practical application.
Adaptability to changing environments is another key concern. Many navigation systems rely on pre-mapped layouts or presume static surroundings, rendering them ineffective in real-world urban environments with ever-changing impediments. Efforts to integrate external infrastructure, such as public surveillance cameras, aim to address this issue, but reliance on fixed installations limits scalability (Jain et al., 2023). Similarly, inertial-based indoor navigation systems exhibit localization drift over time, resulting in positioning errors (Tsai et al., 2024).
Excessive or incorrect information frequently causes confusion and reduces trust in assistive tools (Gonzalez Penuela et al., 2024). AI hallucinations, where generated responses are misleading or overly detailed, present additional concerns, particularly in multimodal accessibility tools (Seo et al., 2024). These inconsistencies diminish the reliability of AI-based assistance, making it difficult for users to rely on such systems in critical situations.
Customization and user flexibility are frequently ignored. Numerous assistive solutions adopt a uniform approach, neglecting the varied needs of users. Although multimodal interfaces have enhanced accessibility for data visualization, they often fall short in providing customization for various degrees of visual impairment (Zhang et al., 2024). Furthermore, studies emphasize the differing information preferences of people who are blind from birth compared to those who lost their sight later, but many navigation technologies do not accommodate these differences (Hoogsteen et al., 2022).
Scalability and inclusivity continue to be significant issues. Educational materials frequently fall short in providing adequate support for BLV learners, necessitating additional assistive tools for fundamental accessibility (Shaheen, 2024). The lack of universally applicable solutions worsens exclusion, highlighting the necessity for more inclusive and scalable technologies.
3. Methodology
This section describes the navigation approaches explored in this study, including different pathfinding schemes, AI-powered scene description, and our proposed image-processing-based technique for BLV individuals’ real-time navigation.
3.1. Navigation Schemes
To guide BLV individuals effectively, we consider multiple navigation schemes that define movement direction and obstacle-free pathfinding:
3.1.1. Left/Forward/Right (DOF 3)
This approach provides three degrees of freedom (DOF), allowing users to navigate in left, forward, or right directions based on obstacle-free regions. It simplifies decision-making but may lack precision in complex environments.
3.1.2. In Degree (DOF 180)
A more granular approach, where the free path is determined across a 180-degree field of view, offering finer directional control for more precise navigation.
3.1.3. Clock Hour Hand (DOF 13)
Inspired by the clock-face representation, this method divides the navigation space into 13 distinct directions, mimicking how blind individuals often conceptualize orientation (e.g., ”move toward 12 o’clock” for straight ahead). This approach balances accuracy and user intuitiveness, making it a practical solution for real-world applications.
Figure 1. Diagram of three different navigation schemes to define movement direction.
[…]
A woman and a girl in white dresses sit in an open car.
Figure 2. Flow diagram of all our approaches, including the proposed one, for identifying the free path for BLV.
[…]
A woman and a girl in white dresses sit in an open car.
3.2. AI-Assisted Navigation Approaches
This study explores six different navigation approaches, five of which leverage AI models for obstacle avoidance and pathfinding. In contrast, the proposed Approach 6 introduces a computationally efficient image processing algorithm using Depth-First Search (DFS) to determine the longest free path. These approaches vary in input data, processing techniques, and model complexity, aiming to balance accuracy, efficiency, and real-time usability.
3.2.1. Approach 1: 2D RGB Image to Vision Language Models (VLMs)
In this approach, we use 2D RGB images with a resolution of 64×64 pixels to simplify processing, reduce computational complexity, and optimize inference time. The images are then passed to Vision Language Models (VLMs), which analyze the scene and infer the best path for navigation based on visual features. While this method benefits from VLMs’ reasoning capabilities, its reliance on RGB data limits depth perception and precise obstacle avoidance.
3.2.2. Approach 2: Depth Images to VLMs
64×64 depth images are used instead of RGB images to enhance depth perception. Depth information allows VLMs to understand spatial relationships more effectively, improving obstacle detection and navigation decisions. VLMs still require significant computational power for real-time processing, making this approach less efficient for low-resource devices.
3.2.3. Approach 3: Combined RGB and Depth Images to VLMs
This approach combines 64×64 RGB and depth images, providing VLMs with visual and depth information for pathfinding. The model gains a richer understanding of the environment by fusing these two modalities, leading to improved accuracy. Consequently, the increased input complexity results in higher computational costs, making real-time implementation more challenging.
3.2.4. Approach 4: Patch-Based Depth Representation to LLMs
Instead of using full-resolution depth images, this method processes depth data in a 5×5 grid of patch values, significantly reducing input size while retaining essential spatial details. The processed depth patches are fed into LLMs, which analyze the segmented depth information for navigation. Although this method lowers computational requirements, the reduction in resolution may impact accuracy in complex environments.
3.2.5. Approach 5: Classification-Based Pathfinding using Vision Transformers (ViT)
This approach formulates pathfinding as a classification task, where a Vision Transformer (ViT) processes images to predict the best navigation direction. ViTs are powerful deep-learning models that capture long-range dependencies within an image. To enable this, we manually annotated a dataset with directional labels and trained a ViT model to learn the mapping between image features and navigation decisions. However, ViTs require extensive computational resources and large-scale training data, limiting their feasibility for real-time deployment in assistive navigation applications.
3.3. Approach 6: DFS-Based Pathfinding from Patched Depth Images (Proposed)
While AI-assisted solutions have demonstrated strong performance in complex perception tasks, they often suffer from significant limitations in real-world applications. These deep learning-based approaches typically demand substantial computational resources, rely on high-end hardware, and consume considerable energy—making them inefficient and unsuitable for real-time or on-device use, particularly in low-power or embedded systems. Additionally, they often require large annotated datasets for training and may struggle to generalize well across diverse, unseen environments.
To address these challenges, we propose a computationally efficient image processing algorithm for pathfinding. Our method leverages depth images obtained from monocular depth estimation and utilizes a 15×15 patch-based depth representation. Instead of relying on deep learning, we implement a Depth-First Search (DFS) algorithm to traverse the depth image and identify the longest free path, ensuring a clear and safe route for navigation. As illustrated in Figure 4, this approach offers several advantages, including lightweight computation, real-time performance, and greater interpretability.
The algorithm is designed based on two key assumptions:
  • First, the depth scene generally transitions from light to dark, meaning the algorithm will prioritize paths that follow a decreasing intensity trend.
  • Second, the search starts from the lower bottom middle of the image, as this corresponds to the natural position of a phone held in front of a person.
The search progresses upwards, exploring valid paths while adhering to these constraints. The diagram (Figure 3) provides a detailed representation of the algorithm.
DFS-Based Optimal Pathfinding Algorithm:
FUNCTION PathFinder(matrix, start_row, start_col, threshold)
DEFINE recursive DFS to explore paths:
ADD current position to the path
COMPUTE avg_intensity of six upper pixels:
       (up-left, up-up-left, up, up-up, up-right, up-up-right)
IF avg_intensity $<$ threshold OR no valid moves THEN
ADD path to results and RETURN
IF top rows reached OR intensity diff. $>$ threshold THEN
ADD path to results and RETURN
FOR EACH valid neighbor (up, up-left, up-right) DO
IF neighbor value $\leq$ current value THEN
RECURSE DFS with updated path
IF no further moves THEN
ADD path to results
CALL DFS from (start_row, start_col)
RETURN the longest and most straight-forward path
Figure 3. Diagram for DFS-based pathfinding algorithm illustrating path exploration and optimal route selection.
4. Results and Discussion
Table 2. Performance Comparison for Indoor Scenes
Approaches Model Tested on 300 Images Tested on 30 Images
MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$ Avg. Response Time (s) $\downarrow$ MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$ Avg. Response Time (s) $\downarrow$
Approach 1 (Real(64x64) $\rightarrow$ VLM) gemini-1.5-pro 49.98 15.67 4.306 47.28 15.44 4.324
BLIP-2 56.02 10.08 3.288 55.50 12.00 3.260
LLaVA-1.5 58.78 9.44 3.040 58.26 9.98 3.110
Approach 2 (Depth(64x64) $\rightarrow$ VLM) gemini-1.5-pro 46.04 15.45 4.776 44.66 16.92 4.825
BLIP-2 48.00 13.98 3.562 47.78 14.67 3.616
LLaVA-1.5 49.32 12.67 3.041 50.02 10.78 3.120
Approach 3 (Real+Depth $\rightarrow$ VLM) gemini-1.5-pro 42.78 16.04 5.118 42.00 16.67 5.161
BLIP-2 47.24 14.32 3.340 46.50 14.75 3.360
LLaVA-1.5 48.91 11.08 3.016 48.77 10.62 2.997
Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 47.90 18.33 3.257 40.00 32.38 3.257
Phi-3-mini-4k-instruct 54.35 14.00 3.094 51.00 20.00 3.094
gemini-1.5-pro 39.05 20.67 5.750 37.50 30.00 5.750
Approach 5 (Classification $\rightarrow$ ViT) Train270_Test30 ViT-Base 52.50 13.29 0.087
Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS)(Proposed) Pathfinder Algorithm 39.81 32.04 0.377 25.78 37.67 0.399
Table 3. Performance Comparison for Outdoor Scenes
Approaches Model Tested on 300 Images Tested on 30 Images
MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$ Avg. Response Time (s) $\downarrow$ MAE (degree) $\downarrow$ Accuracy (%) $\uparrow$ Avg. Response Time (s) $\downarrow$
Approach 1 (Real(64x64) $\rightarrow$ VLM) gemini-1.5-pro 39.94 16.04 4.45 40.50 15.00 4.4497
BLIP-2 40.67 14.72 3.65 42.00 14.00 3.65
LLaVA-1.5 47.00 10.12 3.288 48.50 8.66 3.288
Approach 2 (Depth(64x64) $\rightarrow$ VLM) gemini-1.5-pro 39.18 15.5 5.1912 40.00 13.33 5.1912
BLIP-2 41.24 11.67 3.69 41.67 11.14 3.69
LLaVA-1.5 47.58 10.38 3.32 47.62 9.95 3.32
Approach 3 (Real(64x64) + Depth(64x64) $\rightarrow$ VLM) gemini-1.5-pro 40.50 14.66 5.109 41.74 13.04 5.109
BLIP-2 42.39 14.00 3.557 44.00 12.78 3.557
LLaVA-1.5 49.45 11.32 3.42 52.46 7.66 3.42
Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) Phi-3.5-mini-instruct 40.50 15.33 3.72 42.00 13.33 3.72
Phi-3-mini-4k-instruct 49.05 10.00 3.635 54.00 6.67 3.635
gemini-1.5-pro 40.13 16.05 5.615 41.00 16.67 5.615
Approach 5 (Classification $\rightarrow$ ViT) (Train270_Test30) ViT-Base 41.38 13.79 0.091
Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS)(Proposed) Pathfinder Algorithm 36.45 34.67 1.136 26.00 38.33 1.332
4.1. Evaluation Metrics
To evaluate the performance of the proposed directional prediction algorithm, we consider three key metrics: Mean Absolute Error (MAE), Accuracy, and Average Response Time.
4.1.1. Mean Absolute Error (MAE)
MAE measures the average deviation between the predicted navigational direction and the ground truth, both expressed in clock-hour format. The ground truth directions were manually annotated separately for 300 images from the NYU Depth V2 (Nathan Silberman and Fergus, 2012) dataset (indoor scenes) and 300 images from the DIODE (Vasiljevic et al., 2019) dataset (outdoor scenes). For example, if the ground truth direction is annotated as 11:00 (i.e., $330^{\circ}$), and the predicted direction is 10:00 (i.e., $300^{\circ}$), then the angular deviation is $|330^{\circ}-300^{\circ}|=30^{\circ}$, which corresponds to a 1-hour error. Formally, MAE is calculated as:
(1) $\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|\theta_{i}^{\text{pred}}-\theta_{i}^{\text{gt}}\right|$
where $N$ is the total number of images, $\theta_{i}^{\text{pred}}$ is the predicted direction in degrees for the $i^{\text{th}}$ image, and $\theta_{i}^{\text{gt}}$ is the corresponding ground truth direction.
4.1.2. Accuracy
Accuracy is defined as the proportion of predictions that exactly match the ground truth in clock-hour format. That is, if both predicted and annotated directions are the same (e.g., both are 2:00), it is considered a correct prediction.
(2) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\theta_{i}^{\text{pred}},\theta_{i}^{\text{gt}})$
where the indicator function $\delta(a,b)$ is defined as:
(3) $\delta(a,b)=\begin{cases}1,&\text{if }a=b\\ 0,&\text{otherwise}\end{cases}$
4.1.3. Average Response Time
The average response time refers to the mean computational time (in milliseconds) taken by the algorithm to generate the predicted direction for each input image. This metric reflects the system’s suitability for real-time applications.
Among the metrics, MAE is the most crucial as it directly quantifies angular deviation from the ground truth and thus determines the reliability of navigation. While high Accuracy ensures consistency, a low MAE guarantees that even near-misses are within acceptable tolerance levels. Response Time, though secondary, is significant for real-time deployment, particularly in assistive navigation applications for visually impaired users.
Figure 4. An in-depth visualization of our proposed method for identifying the optimal path using the Pathfinder algorithm, applied to a sample input from the DIODE (Vasiljevic et al., 2019) dataset.
[…]
A woman and a girl in white dresses sit in an open car.
4.2. Indoor Scene Performance
The performance comparison for indoor scenes, as shown in Table 2, reveals key insights into the effectiveness and efficiency of the proposed and baseline approaches. Notably, the proposed Approach 6 (Depth(15x15(patch_value)) $\rightarrow$ DFS) demonstrates competitive performance across key metrics, balancing mean absolute error (MAE), accuracy, and average response time.
For the indoor scene with 30 test images, Approach 6 achieves an MAE of 25.78 degrees with an accuracy of 37.67% and a notably fast average response time of 0.399 seconds. While the MAE value is slightly better than some LLM-based approaches (e.g., Approach 4 (Depth(5x5(patch_value)) $\rightarrow$ LLM) with 37.5 degrees MAE using gemini-1.5-pro), the proposed method significantly outperforms others in terms of computational efficiency. The proposed approach achieves an average response time of 0.399 seconds, which is significantly faster than VLM and LLM-based methods, where response times range between 3.094 and 5.75 seconds for other approaches. This indicates that the DFS-based Pathfinder Algorithm is highly optimized for rapid processing, making it a suitable choice for real-time navigation applications.
Moreover, the classification-based Approach 5 (ViT-Base) exhibits the lowest response time of 0.087 seconds, but its MAE of 52.5 degrees and accuracy of 13.29% indicate a trade-off between speed and prediction quality. This suggests that while ViT-based approaches are computationally efficient, they may lack the precision required for fine-grained obstacle navigation tasks.
4.3. Outdoor Scene Performance
Table 3 presents the performance comparison for outdoor scenes, where the proposed Approach 6 again demonstrates remarkable efficiency and improved accuracy under specific conditions. For the set of 30 test images, Approach 6 achieves an MAE of 26.00 degrees with an accuracy of 38.33%, outperforming all other approaches in both metrics. This represents a significant improvement over the best-performing VLM-based approach (Approach 4 with gemini-1.5-pro), which achieved an MAE of 41.00 degrees and an accuracy of 16.67%.
The efficiency of the proposed approach is further highlighted by its response time of 1.332 seconds, which, while slower than the ViT-based method (0.091 seconds), is still markedly faster than all VLM and LLM-based methods. This suggests that the Pathfinder Algorithm strikes an effective balance between accuracy and computational efficiency, making it well-suited for real-world applications where rapid decision-making is critical.
Interestingly, the proposed Approach 6 also performs consistently across indoor and outdoor environments, with only a slight increase in response time when transitioning to outdoor scenes. This consistency indicates the robustness of the DFS-based method in diverse environments, which is essential for navigation systems assisting blind and low-vision (BLV) individuals.
4.4. Comparative Analysis and Key Observations
  • Accuracy vs. Efficiency Trade-off: While VLM and LLM-based approaches achieve competitive accuracy in some cases, their response times are significantly higher, making them less practical for real-time navigation. The proposed DFS-based method (Approach 6) provides a favorable trade-off, offering a substantial improvement in response time (approximately 3-10 times faster) while maintaining competitive accuracy (around 2-3 times better), without a severe loss in performance. This demonstrates the method’s ability to balance precision with efficiency, making it more suitable for real-time applications.
  • Environment-Specific Performance: The proposed method exhibits superior performance in outdoor scenes, achieving the lowest MAE and highest accuracy among all approaches. In particular, Approach 6 achieves a 106-152% improvement in accuracy and a 16-24% reduction in MAE compared to other methods, suggesting that the algorithm is particularly well-suited for more complex and variable environments like outdoor settings.
  • Model-Specific Observations: Among LLM-based methods, gemini-1.5-pro consistently outperforms other models (Phi-3.5-mini-instruct and Phi-3-mini-4k-instruct) in both accuracy and MAE. However, the increased computational cost associated with these models limits their real-time applicability. In contrast, the proposed DFS-based method achieves a similar or better performance in terms of accuracy, with a significantly lower response time (roughly 3-10 times faster), making it more suitable for practical deployment.
  • Scalability and Real-Time Suitability: The rapid processing times of the proposed method indicate its potential for deployment in real-time assistive navigation systems. Specifically, Approach 6 offers a 3-10 times faster response time than other methods, which is critical for time-sensitive applications. Furthermore, its robustness across both indoor and outdoor scenes suggests scalability to diverse environments, ensuring that the method remains effective in varying real-world conditions.
The DFS-based optimal pathfinding approach offers several key advantages that make it especially suitable for real-time navigation assistance for blind or visually impaired (BLV) individuals. Its lightweight DFS-based pathfinding algorithm demands significantly less computational power than conventional deep learning models, allowing efficient operation on mobile devices without high-end hardware. By utilizing depth variations, the method accurately distinguishes walkable paths from obstacles, enhancing path detection reliability. Additionally, it ensures real-time responsiveness by providing rapid path updates, which supports seamless interaction in dynamic environments. Furthermore, the system is both user-friendly and cost-effective, requiring only a single-camera smartphone and eliminating the need for additional sensors or specialized equipment.
Overall, the proposed Approach 6 emerges as a promising solution for assistive navigation, offering a balanced trade-off between accuracy and computational efficiency. The method stands out, particularly in outdoor environments where rapid decision-making is crucial, and it consistently delivers superior results compared to the alternatives.
5. Usability Study
To assess the usability of PathFinder and the Topline System, we conducted a study with 15 blind or visually impaired (BLV) participants from diverse age groups, education levels, mobility aids, and geographical locations. The study sought to answer critical questions such as:
  • How well does the established system meet the demands of BLV individuals, and how do they rate its effectiveness?
  • How do alternative navigation schemes perform in terms of usability and accuracy?
  • What were the biggest issues that participants encountered when using the app?
  • What improvements or feature requests did the participants make?
Participants were chosen based on various blindness characteristics, such as congenital blindness, visual loss caused by medical disorders (diabetic retinopathy, glaucoma, cataract, etc.), and accidents. The sample also varied in age, education level, mobility aids, and technology knowledge, allowing for a more thorough review of the system. Table 7 provides detailed participant information, including their blindness profile, approach procedures, ease of use ratings, response times, and method rankings. Additionally, participants provided feedback on indoor (7 participants) and outdoor (8 participants) navigation experiences, using a five-point Likert scale, with results visualized through boxplots showing percentile distributions.
5.1. Participant Demographics and Mobility Aids
The study included 15 participants, aged 25 to 80 years, with varying causes of blindness, including birth conditions, accidents, diabetic retinopathy, glaucoma, strokes, cataracts, and macular degeneration. Educational backgrounds ranged from no formal education to university graduates, and participants were from both urban and rural areas across Bangladesh.
Participants used a variety of mobility aids, with the white cane (7 participants) being the most common, followed by smart canes (3), guide dogs (2), GPS-based smart assistants (2), and haptic feedback devices (2). These mobility aids influenced how participants interacted with the tested navigation systems and their expectations for usability and feedback mechanisms.
Table 4 summarizes the participants’ information, including their age, source of blindness, education level, mobility aid utilized, and location.
Table 4. Summary of Survey Participants
Factor Details
Total Participants 15
Age Range 25 - 80 years
Gender Distribution 10 Male, 5 Female
Cause of Vision Loss Birth (2), Accident (2), Glaucoma (1), Diabetic Retinopathy (2), Stroke (2), Cataracts (2), Macular Degeneration (1), Retinitis Pigmentosa (1), Vitamin A Deficiency (1), Age-Related Vision Loss (1)
Mobility Aids White Cane (6), Smart Cane (3), Haptic Feedback Devices (2), Guide Dog (2), GPS-Based Smart Assistants (2)
Testing Environment Indoor (7), Outdoor (8)
Average Duration of App Use 10.13 minutes
5.2. Participant Engagement and Feedback Collection
To ensure a diverse evaluation context, we implemented various engagement strategies during the usability study of our application. Some engaged in conversations with hospital patients, shopkeepers, neighbors, and community members, while others requested feedback from teachers, friends, and family members. This approach provided insights into real-world navigation challenges and user preferences across different social settings. All participants were informed of the purpose of the testing and provided their consent to participate, ensuring that they understood their role in the evaluation process and the significance of their feedback.
5.3. Ease of Use and Response Time
Participants rated the ease of use of Approach 6 on a qualitative scale (Good, Average, Bad):
  • 9 participants rated the system as ”Good”.
  • 4 participants rated it as ”Average”.
  • 2 participants rated it as ”Bad”.
Figure 5. Distribution of Participants’ Ease of Use Ratings.
Figure 6. Distribution of Participants’ Response Time for Providing Feedback on the App.
The response time for participants to complete navigation tasks ranged from 2 to 6 minutes, with an average of approximately 3.8 minutes. Participants using haptic feedback devices or GPS-based smart assistants generally took longer, whereas those using white canes and guide dogs exhibited more varied response times.
5.4. Participant Feedback on Indoor and Outdoor Navigation
Participants evaluated their navigation experience in indoor (7 participants) and outdoor (8 participants) settings using a five-point Likert scale, where 1 = Very Difficult and 5 = Very Easy. The responses were visualized through boxplots in figure 10, illustrating percentile distributions.
5.4.1. Indoor Navigation
Participants generally found indoor navigation more challenging, citing difficulties in identifying obstacles, turns, and structural variations.
5.4.2. Outdoor Navigation
Outdoor navigation received higher Likert ratings, as participants benefited from open spaces, familiar routes, and natural audio cues. However, a few participants rated outdoor navigation due to traffic noise, moving obstacles, and surface irregularities.
Figure 7. Usability ratings (1–5) for navigation schemes across indoor (purple) and outdoor (grey) scenes. Narrow boxplots show scene-wise rating distributions per scheme.
5.5. User Preferences for Navigation Schemes
To determine which navigation strategy felt most intuitive and practical for BLV users, participants ranked three distinct approaches: Left/Forward/Right (DOF 3), Clock Hour Hand (DOF 13), and In Degree (DOF 180). Participants rated these systems based on their ease of use, accuracy, and real-world application. Ranking results are summarized as follows:
  • Clock Hour Hand scheme was selected as the first choice by 7 out of 15 participants, making it the most preferred option. Users praised its systematic approach to navigation, claiming that it closely corresponds with how they currently understand directions.
  • Left/Forward/Right was ranked first by six participants, particularly those with minimal formal education or limited technological knowledge. They found it more intuitive since it resembled how a sighted guide would advise them. However, in crowded outdoor or cramped indoor environments, they did not prefer this method for long-term use.
  • In Degree was the least popular option, with most participants claiming that precise angles were difficult to understand in real time. However, several individuals with technical skills or prior familiarity with digital navigation found it useful.
Figure 8. Bar chart showing the number of participants who selected each navigation scheme as their first priority, categorized by indoor and outdoor scenes.
5.6. Method Ranking and Participant Preferences
Participants ranked navigation methods based on usability and personal preference. Method 6 received the highest ranking across most participants, indicating its perceived effectiveness. Participants using haptic feedback devices and smart canes showed a preference for Method 5 and 6, while those using white canes had mixed preferences. Participants who rated ease of use as ”Bad” had a more varied ranking of methods, suggesting difficulties in adapting to certain navigation approaches.
5.6.1. User Insights on Methods
Participants provided qualitative feedback on why they appreciated or disliked specific strategies. Why Participants Preferred Method 6 (DFS-Based):
  • ”This way seems the most natural and accurate. I don’t have to stop and think; I simply move.”
  • ”It gives me clear paths quickly without making me confused.”
  • ”Works well in crowded places where I need fast, precise decisions.”
Figure 9. Bar chart showing the number of participants who selected each of the six approaches as their top preference for both indoor and outdoor scenes.
Figure 10. Rankings (1–6) of approaches across participants, with narrow boxplots showing rank distributions for each approach, where lower ranks indicate higher preference.
5.7. Challenges with Other Approaches
The table 5 presents the challenges faced by all participants across various free pathfinding approaches.
Table 5. User challenges associated with various pathfinding methods.
Approach Challenges
Approach 1 & 2 Too slow, sometimes the app took too long to respond. There was a delay between asking and getting a response, which made me hesitate. Good explanations, but the lag made it hard to use in real-time.
Approach 3 & 4 The app gave good directions, but sometimes it was unclear if I should turn or move forward. I liked the visual cues, but they were confusing in tight spaces. Works well when there’s enough lighting, but struggles in dim areas.
Approach 5 (ViT-Based Classification) Fast but not always accurate, which is risky in unfamiliar areas. Quick to react, but occasionally misinterpreted obstacles. It missed a few important signs, which made me nervous while navigating.
5.8. Learning the App & Adaptability
One of the most important components of usability is how rapidly BLV users learn and adapt to the application. To evaluate this, participants provided input on Time Taken to Learn the App (How long it took them to feel comfortable using it), Training Needed (Whether they required assistance or self-learning was adequate), Feature Upgrade Requests (Any suggested enhancements for greater usability). Key Insights:
  • Table 6 demonstrates that 66.67% of participants learnt the app in less than a minute, demonstrating a straightforward design.
  • Users Who Did Not Need Training: These participants, particularly those who were familiar with assistive technology, considered the app easy to use.
  • Users Who Needed Training: Some users (particularly elderly ones) preferred hands-on instruction before feeling comfortable using the program.
Table 6. Learning Time Distribution of Participants
Learning Time Number of Participants Percentage
$\leq$ 30 seconds 1 6.67%
30-60 seconds 9 60%
More than 1 min 5 33.33%
5.9. Key Findings from the Usability Study
Based on user input and review, several major insights arose about the app’s usability and pathfinding methods for BLV persons.
5.9.1. What Worked Well
  • Diverse navigation needs: Mobility aids and environmental factors significantly influenced participant preferences.
  • Indoor navigation challenges: Indoor environments presented more difficulties, reflected in lower Likert ratings and higher response times.
  • Outdoor navigation advantages: Participants found outdoor navigation easier, though some faced challenges with environmental distractions.
  • Method 6 (DFS-Based) was the clear favorite: 80% of participants (12 out of 15 from table 7) selecting it as their top priority due to various advantages such as good accuracy, speed, and ease of use.
5.9.2. Challenges Identified
  • Some users had difficulty interpreting methods: A few participants found certain navigation strategies, particularly In Degree (DOF 180), to be very complex.
  • 33% of participants required training, with the majority being older or unfamiliar with technology.
  • Users requested clearer and more frequent voice instructions to improve usability.
  • Response time variability: While the app functioned well in most cases, some users reported slight delays in direction updates when obstacles changed rapidly.
  • Request for haptic feedback: Users suggested incorporating haptic feedback, as voice output may not be practical in noisy or crowded places.
These findings underscore the importance of adaptive navigation solutions that accommodate different environments, mobility aids, and real-world challenges encountered by BLV individuals.
Table 7. Demographic details, mobility aids, approach procedures, ease of use, response times, and method rankings from participant feedback in the usability study.
PID Blindness Profile Age Gender Education Location Mobility Aid Participant Interaction Approach Ease of Use Response Time Method Ranking
1 Blind by birth 25 Male No Formal Education Faridpur White Cane Talk outside the eye hospital in Faridpur Average 2 minutes 6¿3¿5¿4¿2¿1
2 Blind due to accident 47 Female Higher Secondary Dhaka White Cane Request for permission to use the app Average 3 minutes 6¿3¿4¿5¿1¿2
3 Due to Glaucoma 74 Male Retired School Teacher Noakhali White Cane Requested a school teacher’s feedback Good 4 minutes 6¿3¿4¿5¿2¿1
4 Diabetic retinopathy 54 Female Secondary Education Barishal Smart Cane Talked to a stranger at Faridpur Diabetic Hospital Average 4 minutes 6¿5¿3¿4¿2¿1
5 Stroke 69 Male Retired Doctor Khulna Haptic Feedback Devices Requested permission to use the app Good 2 minutes 6¿5¿3¿4¿1¿2
6 Vitamin A deficiency 62 Male Former NGO Worker Faridpur Smart Cane Collected feedback from a passerby on Newmarket Street Good 5 minutes 5¿6¿4¿3¿2¿1
7 Stroke 71 Female Retired Lawyer Faridpur White Cane Collected feedback from a patient at BSMRH Hospital Bad 4 minutes 6¿3¿4¿1¿2¿5
8 Macular degeneration 80 Male Retired College Professor Jessore GPS-Based Smart Assistants Asked a neighbor for feedback Average 2 minutes 3¿5¿6¿1¿2¿4
9 Blind due to accident 48 Male Vocational Education Jessore Guide Dog Asked a friend for feedback Good 6 minutes 6¿3¿5¿4¿2¿1
10 Cataract (Untreated) 65 Male Primary Education Chattogram White Cane Talked to a shopkeeper near his house Good 3 minutes 6¿3¿4¿5¿1¿2
11 Diabetic Retinopathy 52 Female Secondary Education Sylhet GPS-Based Smart Assistants Discussed in a community center Bad 5 minutes 5¿6¿3¿4¿2¿1
12 Cataracts 70 Male Higher Secondary Rajshahi White Cane Spoke with a friend at the bus station Good 4 minutes 6¿5¿4¿3¿2¿1
13 Retinitis Pigmentosa 35 Male Bachelor’s Degree Rangpur Smart Cane Collected feedback at a local café Good 3 minutes 6¿4¿3¿5¿1¿2
14 Birth Condition 29 Female University Graduate Mymensingh Guide Dog Asked a university friend for feedback Good 2 minutes 6¿5¿3¿4¿2¿1
15 Age-related vision loss 76 Male Secondary School Mymensingh Haptic Feedback Devices Tested the app while walking with a family member Good 6 minutes 6¿3¿4¿1¿2¿5
Table 8. Participant demographics, navigation scheme preferences, and user feedback on the usability of the app.
PID Vision Ability Scene Type Time to Learn App Duration of Use Navigation Scheme Ranking User Feedback
1 Totally blind Outdoor 80 secs 6 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree This method is well-suited for complex scenarios while maintaining a lower inference time. The 6th method is the most optimal choice for low latency.
2 Light perception only Indoor 40 secs 10 minutes Clock Hour Hand ¿ Left/Forward/Right ¿ In Degree I think this method is tailored for outdoor scenarios, ensuring efficient performance in dynamic environments.
3 Tunnel vision Indoor 50 secs 12 minutes Clock Hour Hand ¿ Left/Forward/Right ¿ In Degree I feel like the method which took the least amount of time is perfect. The 6th method is my choice for low latency.
4 Blurry vision Outdoor 70 secs 8 minutes Clock Hour Hand ¿ Left/Forward/Right ¿ In Degree The clock hour-hand scheme is an effective way to represent direction. The 6th method stands out for its efficiency.
5 Partial vision loss Indoor 45 secs 12 minutes Clock Hour Hand ¿ In Degree ¿ Left/Forward/Right The clock hour-hand scheme is more intuitive, but the 6th method balances efficiency and accuracy.
6 Difficulty seeing at night Outdoor 30 secs 10 minutes In Degree ¿ Clock Hour Hand ¿ Left/Forward/Right Method 5 has the lowest inference time, ideal for fast applications. Method 6 is more balanced but not the fastest.
7 Partial vision loss Outdoor 40 secs 9 minutes Clock Hour Hand ¿ Left/Forward/Right ¿ In Degree Method 6 delivers better results in most cases. I appreciate its way of representing direction.
8 Distorted vision Indoor 55 secs 10 minutes Clock Hour Hand ¿ In Degree ¿ Left/Forward/Right I prefer method 3 as it provides better results, and using the clock hour-hand direction is a good choice.
9 Light perception only Indoor 50 secs 14 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree The left/forward/right scheme is easier to understand. Approach 6 is innovative and provides a balanced result.
10 Totally blind Outdoor 75 secs 10 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree The left/forward/right method feels natural, like how someone would guide me in real life. Method 6 worked really well.
11 Blurry vision Outdoor 65 secs 12 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree The left/forward/right scheme makes the most sense to me, but the clock hour-hand system sometimes requires more thought.
12 Partial vision loss Outdoor 45 secs 9 minutes Clock Hour Hand ¿ Left/Forward/Right ¿ In Degree The clock hour-hand method was the easiest to follow in real-life situations, providing fast and clear directions.
13 Tunnel Vision Outdoor 50 secs 8 minutes In Degree ¿ Clock Hour Hand ¿ Left/Forward/Right I liked method 6 because it felt natural. The degree-based method was interesting but sometimes required stopping to process.
14 Totally Blind Indoor 70 secs 10 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree I love how the app gives clear instructions. Method 6 was quick, clear, and effortless to understand.
15 Blurry Vision Indoor 50 secs 12 minutes Left/Forward/Right ¿ Clock Hour Hand ¿ In Degree The left/forward/right method was the easiest because it felt natural. Method 6 provided clear instructions and confidence.
Figure 11. Feedback from a participatory study with 15 BLV patients, where responses on indoor (7 participants) and outdoor (8 participants) scenes were rated on a five-point Likert scale and visualized using boxplots to show percentile distributions.
6. Limitations and Potential Future Work
While our suggested navigation system provides an efficient and cost-effective solution for blind and low-vision (BLV) users, some limitations must be addressed in future study.
One significant restriction is that our system does not direct people to their eventual destination, but rather offers a short, obstacle-free path in their immediate surroundings. While this enables safe navigation in local surroundings, incorporating a global path-planning mechanism may improve long-distance mobility and route optimization.
Another limitation is working in low-light circumstances, as the system is based on image-based depth estimate. Depth accuracy diminishes in dark situations, making obstacle detection less accurate. Future enhancements could include the use of infrared or LiDAR-based sensors to increase performance in low-light conditions.
Our technique doesn’t struggle as much in dynamic situations, where moving objects like pedestrians and automobiles present unforeseen challenges. Currently, the technology is intended for static scenarios, which limits its real-time adaptability. However, using haptic feedback instead of voice feedback might help resolve this issue, enabling more dynamic and responsive interaction. Future editions could investigate the use of real-time object tracking and motion prediction algorithms to deal with dynamic impediments more efficiently.
Furthermore, complex surfaces such as stairs and steep slopes cause difficulties for depth estimation, resulting in less precise path recommendations. To address this issue, the depth estimation model must be refined to better differentiate between navigable and non-navigable surfaces. Using terrain categorization algorithms may improve the system’s capacity to manage such situations. Currently, there is no available dataset for this purpose. We have manually annotated a small dataset for testing, but a larger and more diverse dataset for fine-tuning could significantly improve the results of VLSs (Vision Language Systems).
Finally, participant recruitment for testing remains difficult, especially in Bangladesh, where many BLV users are unfamiliar with technology navigation aids. The limited availability to experienced users has an impact on the review process. Future development should concentrate on awareness campaigns and collaborations with local organizations to increase user adoption and collect more detailed feedback.
By addressing these limitations, we hope to improve our system’s stability, usability, and real-world application, resulting in more independent mobility for BLV individuals.
7. Conclusion
This study presented a novel, computationally efficient navigation system designed to enhance mobility for Blind and Low Vision (BLV) individuals. By leveraging a depth-based pathfinding algorithm with a Depth-First Search (DFS) approach, we demonstrated a low-cost, real-time solution that accurately detects free paths in both indoor and outdoor environments. Comparative evaluations showed that our proposed system offers a favorable trade-off between accuracy and computational efficiency, making it a viable alternative to existing AI-powered navigation solutions that require high-end hardware.
Our usability study, conducted with BLV participants, highlighted the system’s effectiveness in real-world settings, with the majority of users preferring our approach for its intuitive directional guidance and low inference time. While outdoor navigation yielded higher satisfaction due to fewer obstacles and broader spaces, indoor navigation posed challenges that require further refinement.
Despite the promising results, our study identified several limitations, including challenges in low-light environments, dynamic obstacle handling, and long-distance navigation. Future research should focus on integrating multi-modal sensor inputs, improving adaptability to rapidly changing environments, and refining global path-planning mechanisms to enhance user autonomy. Additionally, addressing accessibility barriers through collaborations with BLV communities will be essential in optimizing the system’s practical adoption.